OpenStudy (anonymous):

find the x-and y- intercepts of each equation -3x+3y=7

OpenStudy (jdoe0001):

to get x-intercept, set y = 0, solve for "x" to get y-intercept, set x = 0, solve for "y"

OpenStudy (anonymous):

yes but -3 times 0 is 0 but then you have 3/7 so do you divide that and keep it or do you not divide this answer and leave 3/7 as a x y to graph

OpenStudy (anonymous):

\[-3x+3y=7\] \[-3x+3(0)=7\] \[\frac{3 }{3}y = \frac{ 7 }{ 3} ?\]

OpenStudy (anonymous):

yes but -3 times 0 is 0 but then you have 3/7 so do you divide that and keep it or do you not divide this answer and leave 3/7 as a x y to graph

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i dont know if i should divide -3/7 to a decimal or just leave it alone

OpenStudy (anonymous):

u can leave it alone. unless they specifically ask u to to put it in decimal form

OpenStudy (anonymous):

no they just ask for the x and y but how will i be able to graph a fraction unless it is a slope

OpenStudy (anonymous):

i got \[(-\frac{ 7 }{ 3 },0)\] for x and \[y= (0, \frac{ 7 }{ 3})\]